OpenStudy (anonymous):

WHAT ABOUT THIS ONE? How much should be deposited in an account paying 6% interest compounded monthly in order to have a balance of $2000 after 2 years?

OpenStudy (anonymous):

2000 = p (1+(0.06/12)) ^ (24) Solve for p.

OpenStudy (anonymous):

Using Future Value formula FV = PMT * ( ( (1 + i)^n - 1) / i ) FV = Future Value = 2000 PMT = Periodic Payment Amount <= we look for this i = interest rate per period = 6% = 0.06 n = number of periods= 2 years x 12 months = 24 2000 = PMT * (((1 + 0.06)^24 -1)/ 0.06) = PMT * 50.82 PMT= 2000/ 50.82 = 39.35 dollars/ month

OpenStudy (anonymous):

I think that he is looking for the principal on a one time investment, not a growing balance.

OpenStudy (anonymous):

Oops...misunderstanding in info...nevermind...so dfyodor formula should be right...you can keep mine as reference in the type of question they asking you to deposit monthly...